Given the following situations, write A, if it is two-tailed, B if one-tailed (left-


tailed), and C if one-tailed (right-tailed). (10 points)



1. Research on processed foods lessens the life span of a person.


2. A government official claims that the dropout rate for local schools is 25%.


3. A civil engineer is trying to prove that his new method of laying blocks is more


effective than the traditional method.


4. A new stockbroker, named XYZ, claims that their brokerage fees are lower than


that of the current fee.


5. A government official claims that the dropout rate for local schools is 25%.

1
Expert's answer
2022-05-25T12:56:30-0400

1.

"H_0: \mu\ge\mu_0"

"H_a: \mu<\mu_0"

B one-tailed (left-tailed)


2.

"H_0: \mu=\mu_0"

"H_a: \mu\not=\mu_0"

A two-tailed


3.

"H_0: \mu\le\mu_0"

"H_a: \mu>\mu_0"

C one-tailed (right-tailed)


4.

"H_0: \mu\ge\mu_0"

"H_a: \mu<\mu_0"

B one-tailed (left-tailed)


5.

"H_0: \mu=\mu_0"

"H_a: \mu\not=\mu_0"

A two-tailed